40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14033     Entry time: Fri Jun 29 18:16:32 2018
Author: Jon 
Type: Configuration 
Category: PSL 
Subject: Changes to AUX Optical Layout on PSL Table 

In order to use the 0th-order deflection beam from the AOM for cavity mode scans, I've coaligned this beam to the existing mode-matching/launch optics set up for the 1st-order beam.

Instead of being dumped, the 0th-order beam is now steered by two 45-degree mirrors into the existing beam path. The second mirror is on a flip mount so that we can quickly switch between 0th-order/1st-order injections. None of the existing optics were touched, so the 1st-order beam alignment should still be undisturbed.

Currently the 0th-order beam is being injected into the IFO. After attenuating so as to not exceed 100 mW incident on the fiber, approximately 50 mW of power reaches the AS table. That coupling efficiency is similar to what we have with the 1st-order beam. With the Y-arm cavity locked and the AUX PLL locked at RF offset = 47.60 MHz (an Y-arm FSR), I observed a -50 dBm beat note at Y-end transmission.

Attachment 1: PSL_AUX_SETUP_CHANGE.pdf  408 kB  Uploaded Fri Jun 29 19:17:20 2018  | Hide | Hide all
PSL_AUX_SETUP_CHANGE.pdf
ELOG V3.1.3-